# Constraining the Stellar-to-Halo Mass Relation with Galaxy Clustering and Weak Lensing from DES Year 3 Data

### v2

@@ Abstract
We develop a framework to study the relation between the stellar mass of a galaxy and the total mass of its host dark matter halo using galaxy clustering and galaxy-galaxy lensing measurements. We model a wide range of scales, roughly from $\sim 100\;{\rm kpc}$ to $\sim 100\;{\rm Mpc}$, using a theoretical framework based on the Halo Occupation Distribution and data from Year 3 of the Dark Energy Survey (DES) dataset. The new advances of this work include: 1) the generation and validation of a new stellar mass-selected galaxy sample in the range of $\log M_{\star}/M_{\odot}\sim 9.6$ to $\sim 11.5$; 2) the joint-modeling framework of galaxy clustering and galaxy-galaxy lensing that is able to describe our stellar mass-selected sample deep into the 1-halo regime; and 3) stellar-to-halo mass relation (SHMR) constraints from this dataset. In general, our SHMR constraints agree well with existing literature with various weak lensing measurements. We constrain the free parameters in the SHMR functional form $\log M_{\star}(M_{h})=\log(\epsilon M_{1})+f\left[\log\left(M_{h}/M_{1}\right)\right]-f(0)$, with $f(x)\equiv-\log(10^{\alpha x}+1)+\delta[\log(1+\exp(x))]^{\gamma}/[1+\exp(10^{-x})]$, to be $\log M_{1}=11.559^{+0.334}_{-0.415}$, $\log\epsilon=-1.689^{+0.333}_{-0.220}$, $\alpha=-1.637^{+0.107}_{-0.096}$, $\gamma=0.588^{+0.265}_{-0.220}$ and $\delta=4.227^{+2.223}_{-1.776}$. The inferred average satellite fraction is within $\sim 5-35\%$ for our fiducial results and we do not see any clear trends with redshift or stellar mass. Furthermore, we find that the inferred average galaxy bias values follow the generally expected trends with stellar mass and redshift. Our study is the first SHMR in DES in this mass range, and we expect the stellar mass sample to be of general interest for other science cases.
[…]
@@ 1 Introduction
Understanding the connection between galaxies and dark matter is one of the essential ingredients needed to form a comprehensive interpretation of the observed Universe. Analyses of Large-Scale Structure (LSS) from galaxy surveys can be biased if details of the galaxy-halo connection are ignored (McDonald & Roy, 2009; Baldauf et al., 2012; Krause et al., 2017). In modern analyses, one common approach to side-step this complication is to remove the data points on the small scales and only use data in the linear to quasilinear regime, where a simple prescription of the galaxy-halo connection such as linear galaxy bias is sufficient (see e.g. DES Collaboration, 2021). In some cases, more complicated galaxy bias models on small scales (such as Heymans et al., 2021) are used and marginalize over the model parameters. For either approach, a data-driven model of the galaxy-halo connection on scales below a few Mpc could allow us to significantly improve the cosmological constraints achievable by a given dataset (e.g. Kwan et al., 2016; Mandelbaum et al., 2013; Cacciato et al., 2013; Park et al., 2016; Krause et al., 2017; Singh et al., 2019). On the other hand, understanding the connection between different galaxy samples and their host halos also has implications for galaxy evolution (see Wechsler & Tinker, 2018, for a review of studies for galaxy-halo connection).
@@
In this work we study the combination of two probes for the galaxy-halo connection: galaxy-galaxy lensing and galaxy clustering. Galaxy-galaxy lensing (Tyson et al., 1984; McKay et al., 2001; Hoekstra et al., 2004; Sheldon et al., 2004; Mandelbaum et al., 2005, 2006; Seljak et al., 2005; Johnston et al., 2007; Heymans et al., 2006; Leauthaud et al., 2011) refers to the measurement of the cross-correlation between the positions of foreground galaxies and shapes of background galaxies. Due to gravitational lensing, the images of background galaxies appear distorted due to the deflection of light as it passes by foreground galaxies and the dark matter halos they reside in. As a result, this measurement maps the average mass profile of the dark matter halos hosting the foreground galaxy sample – a direct measurement of the galaxy-halo connection.
@@
Galaxy clustering (Kaiser, 1984; Fry & Gaztanaga, 1993; Leauthaud et al., 2011; Abbott et al., 2018) refers to the measurement of spatial distribution of galaxies. If we know how galaxies and dark matter halos are connected, the galaxy distribution can be straightforwardly translated to the distribution of dark matter, which makes them a direct probe of LSS. However, without knowledge of the galaxy-halo connection, galaxy clustering on its own suffers from the degeneracy between galaxy bias and the amplitude of the matter power spectrum. As a result, to extract the information in the galaxy clustering measurements, one has to combine it with other probes such as galaxy-galaxy lensing, in order to yield meaningful constraints on the galaxy-halo connection and cosmology (van den Bosch et al., 2003; Tinker et al., 2005; Seljak et al., 2005). Combining the two has been proved (Yoo et al., 2006; Cacciato et al., 2009) to be sensitive to cosmological information, and particularly to $\sigma_{8}$ and $\Omega_{m}$, since this combination probes the shape and amplitude of the halo mass function on small scales and the overall matter density and the bias of the galaxy sample at large scales.
@@
A common approach to modeling both of these measurements is via the Halo Model (Peacock & Smith, 2000; Seljak, 2000; Ma & Fry, 2000; Cooray & Sheth, 2002) and the Halo Occupation Distribution (HOD) framework (Peacock & Smith, 2000; Scoccimarro et al., 2001; Berlind & Weinberg, 2002; Tinker et al., 2007; Zheng et al., 2007; Zehavi et al., 2011; White et al., 2011), where most of the complexities lie within the HOD framework. The foundational assumption in the basic HOD framework is that a central galaxy is considered to exist in the centers of the host halos and the satellite galaxies are distributed within the halo (Kravtsov et al., 2004). The particular way that the galaxies are distributed depends on the mass of the host halo.
@@
Various alternatives of the HOD framework have been proposed in the literature. For example, the conditional luminosity function $\Phi(L|M_{h})$ provides the average number of galaxies of luminosity $L\pm dL$ that reside in a halo of mass $M_{h}$ (Tasitsiomi et al., 2004; Yang et al., 2004; Van Den Bosch et al., 2004; Vale & Ostriker, 2006; Cooray, 2006); the conditional stellar mass function $\Phi(M_{\star}|M_{h})dM_{\star}$ describes the average number density of galaxies in a halo of mass $M_{h}$ with stellar masses in the range of $M_{\star}\pm dM_{\star}$ (Yang et al., 2007; Moster et al., 2010; Behroozi et al., 2013); and abundance matching relies on simulations and assumes that there is there is a monotonic correspondence between halo mass (or circular velocity) and galaxy stellar mass (or luminosity) (e.g., Tasitsiomi et al., 2004; Conroy et al., 2006; Kravtsov et al., 2004; Vale & Ostriker, 2006; Conroy & Wechsler, 2009; Drory et al., 2009; Moster et al., 2010; Behroozi et al., 2013; Guo et al., 2010).
@@
In this paper we present a joint analysis of galaxy-galaxy lensing and galaxy clustering using a new stellar mass-selected galaxy sample from the first three years (Y3) of the Dark Energy Survey (DES) data. We model scales deep in the 1-halo regime using an HOD modeling framework based on our previous work in Zacharegkas et al. (2022).
@@
In Zacharegkas et al. (2022), we analyzed the galaxy-galaxy lensing signal from two lens galaxy samples that were used by DES Y3 for the main cosmological studies (e.g. Abbott et al., 2022; Prat et al., 2022). Our HOD model was widely based on Zheng et al. (2007); Zehavi et al. (2011) and the halo occupation was inferred as a function of the halo mass. The HOD constraints in that work were important in various aspects of understanding the DES Y3 cosmology results, but are not suited for understanding the general galaxy population given that the lens samples had various selections that were imposed to optimize their photometric redshifts. In this work, we expand upon Zacharegkas et al. (2022) in three main ways, which we discuss in what follows. First, we incorporate theory predictions for galaxy clustering on small scales in addition to galaxy-galaxy lensing. In addition, our new model accounts for the stellar mass-information of the lens galaxies. This is done by direct parametrization of the HOD, but is equivalent to using, for example, the conditional stellar mass function mentioned before. Finally, we construct and validate a new, more generic stellar-mass-selected lens sample compared to what was used in Zacharegkas et al. (2022). This allows us to constrain the stellar-to-halo-mass relation of DES galaxies. We also note here that, the new stellar mass-selected galaxy sample can be used in future analyses of other science cases.
@@
Throughout this study we fix the $\Omega_{m}$ and $\sigma_{8}$ cosmological parameters to the best-fit for the MagLim sample from DES Y3 (Abbott et al., 2022), while we use the best-fit cosmology from Planck 2018 (Planck Collaboration, 2020) for the rest of them: $\Omega_{\rm m}=0.339$, $\sigma_{8}=0.733$, $\Omega_{\rm b}=0.0486$, $H_{0}=67.37\;{\rm km/s/Mpc}$, $n_{s}=0.9649$, $\Omega_{\nu}h^{2}=0.0006$, where $h$ is the Hubble constant in units of $100\;{\rm km/s/Mpc}$. Since we consider the $\Lambda$-Cold Dark Matter ($\Lambda$CDM) cosmological model, we set $w=-1$ for the dark energy equation of state parameter. We test in Section 7.2 the sensitivity of our results to the assumed cosmological model. In addition, all the halo masses use the definition of $M_{\rm 200c}$, based on the mass enclosed by radius $R_{\rm 200c}$ so that the mean density of a halo is 200 times the critical density at the redshift of the halo. However, our results are valid for any common choice of mass definition.
@@ 2 The Halo Occupation Distribution model
In this section we discuss how we model the Halo Occupation Distribution (HOD) for this work. This model is based on Zacharegkas et al. (2022) and it is modified so that the stellar-mass information of the galaxies is accounted for. This is done by incorporating a stellar-to-halo-mass relation (SHMR) into the HOD framework. Below we introduce the SHMR used in this study (Section 2.1) and we then proceed to describe our HOD model (Section 2.2).
@@ 2.1 The stellar-to-halo-mass relation
Henceforth, we will be using the notation $f_{\rm SHMR}(M_{h})\equiv M_{\star}(M_{h})$ for the SHMR. There are a large number of variants of this relation used in the literature and, for the most part, they are in good agreement. However, especially at large masses, there can exist significant differences. In this work we follow the SHMR form in Kravtsov et al. (2018), which is the same as in Behroozi et al. (2013). In that paper the authors carefully calibrated the SHMR of high-mass halos and found that the relation in that regime is shallower than the HOD constraints by Leauthaud et al. (2012) and the abundance matching analyses by Moster et al. (2013); Behroozi et al. (2013). The authors suggest that the main reason is that in these previous studies the measured stellar-mass of the central galaxies was underestimated significantly.
@@
In the above expressions, we allow $\{M_{1},\epsilon,\alpha,\gamma,\delta\}$ to vary during our fits to the data. This functional form is a power-law with slope $-\alpha$ at the low-mass limit, $M_{h}\ll M_{1}$, and breaks into a different power-law with slope $\gamma$ for high-mass halos, $M_{h}\gg M_{1}$, at the characteristic halo-mass scale $M_{1}$ with a characteristic stellar-mass-to-halo-mass ratio $\epsilon$. Finally, $\delta$ describes the strength of the power-law in the massive end of the relation.
@@ 2.2 The HOD parametrization
For a galaxy sample selected by a stellar-mass threshold $M_{\star}^{t}$, i.e. $M_{\star}\geq M_{\star}^{t}$, we can write down the HOD for the central galaxy similar to the form presented in Zheng et al. (2007):
@@
To derive the HOD for a central occupation conditioned on a range of stellar masses $[M_{\star,1},M_{\star,2}]$, we simply subtract two thresholded HOD’s from Equation (2) at $M_{\star,1}$ and $M_{\star,2}$:
@@
where $f_{\rm sat}$, $M_{\rm sat}$, $M_{\rm cut}$, $\alpha_{\rm sat}$ and $\kappa$ are free parameters, and we have defined $M_{\min}\equiv f_{\rm SHMR}^{-1}(M_{\star}^{t})$. Specifically, $f_{\rm sat}$ is the equivalent of $f_{\rm cen}$ in the central HOD of Equation (2) for the satellite occupation, $\alpha_{\rm sat}$ controls the power-law behaviour of the satellite occupation at large halo masses, and the stellar-mass dependence enters into the satellite modeling via the parameters $M_{\rm sat}$ and $M_{\rm cut}$ (see, e.g. Leauthaud et al., 2011; Zu & Mandelbaum, 2015), that we model as:
@@
where the new free parameters ($B_{\rm sat}$, $B_{\rm cut}$, $\beta_{\rm sat}$, $\beta_{\rm cut}$) have been introduced to allow the satellites to have a different form of HOD from the centrals. As an additional note, as opposed to Zacharegkas et al. (2022), in this current paper we decouple the HOD of the satellites in Equation (2.2) from the central one to give the model additional freedom in order to fit the more complex data we analyze here.
@@
The above approach effectively assumes a tophat distribution of the stellar mass in the range, and has been shown to be a good model for certain datasets (see e.g. Leauthaud et al., 2011). This is also what we will use as our fiducial model in this work. We describe a more general case in Appendix A, where the stellar mass distribution takes on a more flexible form.
@@
×$\qquad\times\left[2\langle N_{c}(M_{h}|M_{\star}^{t})\rangle+\lambda_{s}(M_{h})\rangle\right]\;.$ (9)
@@ 3 Modeling the observables
Equipped with the new HOD model, the main goal in this section is to combine the HOD model with the Halo Model (Seljak, 2000; Ma & Fry, 2000; Peacock & Smith, 2000) to predict our two observables: galaxy-galaxy lensing (Section 3.1) and galaxy clustering (Section 3.2).
@@
as a function of distance $r$. Following Hayashi & White (2008); Zu & Mandelbaum (2015); Clampitt et al. (2017), we combine the 1-halo and 2-halo term in the 3-dimensional real space, as it is found to more accurately describe observations at the quasi-linear regime. In the above, the 1-halo terms can be further decomposed into the central and satellite terms, as introduced in Kravtsov et al. (2004). We describe how we model each of these terms in $P_{\rm gm}$ and $P_{\rm gg}$ in Sections 3.1 and 3.2, respectively. We also note that each of the terms is a function of redshift ($z$) and scale ($k$). The central and satellite terms refer to the same definitions used in the HOD model in Section 2.2. The 1- and 2-halo terminology refers to whether the two tracers of interest (galaxy or dark matter) live in the same halo or two separate halos. There is a delicate choice to make when combining the 1- and 2-halo terms, which we discuss in Section 3.3. Once each term in Equations (10) and (11) are specified, together with the prescription for combining the two terms, it is projected into 2D (the projected spetra $C_{\rm gm}$ and $C_{\rm gg}$, of $P_{\rm gm}$ and $P_{\rm gg}$ respectively), then Fourier transformed into the real-space estimators that we measure ($\gamma_{t}$ and $w$) in Section 3.4. We describe the approach we take for incorporating the redshift and stellar mass uncertainties in Section 3.5, and finally we discuss any additional modeling components in Section 3.6

### v3

@@ Abstract
We develop a framework to study the relation between the stellar mass of a galaxy and the total mass of its host dark matter halo using galaxy clustering and galaxy-galaxy lensing measurements. We model a wide range of scales, roughly from $\sim 100\;{\rm kpc}$ to $\sim 100\;{\rm Mpc}$, using a theoretical framework based on the Halo Occupation Distribution and data from Year 3 of the Dark Energy Survey (DES) dataset. The new advances of this work include: 1) the generation and validation of a new stellar mass-selected galaxy sample in the range of $\log M_{\star}/M_{\odot}\sim 9.6$ to $\sim 11.5$; 2) the joint-modeling framework of galaxy clustering and galaxy-galaxy lensing that is able to describe our stellar mass-selected sample deep into the 1-halo regime; and 3) stellar-to-halo mass relation (SHMR) constraints from this dataset. In general, our SHMR constraints agree well with existing literature with various weak lensing measurements. We constrain the free parameters in the SHMR functional form $\log M_{\star}(M_{h})=\log(\epsilon M_{1})+f\left[\log\left(M_{h}/M_{1}\right)\right]-f(0)$, with $f(x)\equiv-\log(10^{\alpha x}+1)+\delta[\log(1+\exp(x))]^{\gamma}/[1+\exp(10^{-x})]$, to be $\log M_{1}=11.506^{+0.325}_{-0.404}$, $\log\epsilon=-1.632^{+0.306}_{-0.181}$, $\alpha=-1.638^{+0.108}_{-0.099}$, $\gamma=0.596^{+0.251}_{-0.210}$ and $\delta=3.810^{+2.045}_{-1.811}$. The inferred average satellite fraction is within $\sim 5-35\%$ for our fiducial results and we do not see any clear trends with redshift or stellar mass. Furthermore, we find that the inferred average galaxy bias values follow the generally expected trends with stellar mass and redshift. Our study is the first SHMR in DES in this mass range, and we expect the stellar mass sample to be of general interest for other science cases.
@@ 1 Introduction
Understanding the connection between galaxies and dark matter is one of the essential ingredients needed to form a comprehensive interpretation of the observed Universe. Analyses of Large-Scale Structure (LSS) from galaxy surveys can be biased if details of the galaxy-halo connection are ignored (McDonald2009; Baldauf2012; Krause2017). In modern analyses, one common approach to side-step this complication is to remove the data points on the small scales and only use data in the linear to quasilinear regime, where a simple prescription of the galaxy-halo connection such as linear galaxy bias is sufficient (see e.g. y3-3x2ptkp). In some cases, more complicated galaxy bias models on small scales (such as Heymans2020) are used and marginalize over the model parameters. For either approach, a data-driven model of the galaxy-halo connection on scales below a few Mpc could allow us to significantly improve the cosmological constraints achievable by a given dataset (e.g. Kwan2016; Mandelbaum2013; Cacciato2013; Park2016; Krause2017; Singh2019). On the other hand, understanding the connection between different galaxy samples and their host halos also has implications for galaxy evolution (see Wechsler2018, for a review of studies for galaxy-halo connection).
@@
In this work we study the combination of two probes for the galaxy-halo connection: galaxy-galaxy lensing and galaxy clustering. Galaxy-galaxy lensing (tyson1984; McKay2001; hoekstra2004; Sheldon_2004; mandelbaum2004; Mandelbaum2006; seljak2005; Johnston2007; Heymans2006; leauthaud2011) refers to the measurement of the cross-correlation between the positions of foreground galaxies and shapes of background galaxies. Due to gravitational lensing, the images of background galaxies appear distorted due to the deflection of light as it passes by foreground galaxies and the dark matter halos they reside in. As a result, this measurement maps the average mass profile of the dark matter halos hosting the foreground galaxy sample – a direct measurement of the galaxy-halo connection.
@@
Galaxy clustering (Kaiser1984; Fry1993; leauthaud2011; Abbott2018) refers to the measurement of spatial distribution of galaxies. If we know how galaxies and dark matter halos are connected, the galaxy distribution can be straightforwardly translated to the distribution of dark matter, which makes them a direct probe of LSS. However, without knowledge of the galaxy-halo connection, galaxy clustering on its own suffers from the degeneracy between galaxy bias and the amplitude of the matter power spectrum. As a result, to extract the information in the galaxy clustering measurements, one has to combine it with other probes such as galaxy-galaxy lensing, in order to yield meaningful constraints on the galaxy-halo connection and cosmology (vandenBosch2003; tinker2005; seljak2005). Combining the two has been proved (yoo2006; Cacciato2009) to be sensitive to cosmological information, and particularly to $\sigma_{8}$ and $\Omega_{m}$, since this combination probes the shape and amplitude of the halo mass function on small scales and the overall matter density and the bias of the galaxy sample at large scales.
@@
A common approach to modeling both of these measurements is via the Halo Model (PeacockSmith2000; Seljak2000; Ma2000; cooray2002) and the Halo Occupation Distribution (HOD) framework (PeacockSmith2000; Scoccimarro2001; Berlind2002; Tinker07; zheng2007; zehavi2011; white2011), where most of the complexities lie within the HOD framework. The foundational assumption in the basic HOD framework is that a central galaxy is considered to exist in the centers of the host halos and the satellite galaxies are distributed within the halo (KravtsovHOD2004). The particular way that the galaxies are distributed depends on the mass of the host halo.
@@
Various alternatives of the HOD framework have been proposed in the literature. For example, the conditional luminosity function $\Phi(L|M_{h})$ provides the average number of galaxies of luminosity $L\pm dL$ that reside in a halo of mass $M_{h}$ (Tasitsiomi2004; Yang2004; vandenBosch2004; Vale2006; Cooray2006); the conditional stellar mass function $\Phi(M_{\star}|M_{h})dM_{\star}$ describes the average number density of galaxies in a halo of mass $M_{h}$ with stellar masses in the range of $M_{\star}\pm dM_{\star}$ (Yang2007; Moster2010; Behroozi2013); and abundance matching relies on simulations and assumes that there is there is a monotonic correspondence between halo mass (or circular velocity) and galaxy stellar mass (or luminosity) (e.g., Tasitsiomi2004; Conroy06; KravtsovHOD2004; Vale2006; Conroy2009; Drory2009; Moster2010; Behroozi2013; Guo2010).
@@
In this paper we present a joint analysis of galaxy-galaxy lensing and galaxy clustering using a new stellar mass-selected galaxy sample from the first three years (Y3) of the Dark Energy Survey (DES) data. We model scales deep in the 1-halo regime using an HOD modeling framework based on our previous work in Zacharegkas2022.
@@
In Zacharegkas2022, we analyzed the galaxy-galaxy lensing signal from two lens galaxy samples that were used by DES Y3 for the main cosmological studies (e.g. y3-3x2pt; Prat2022). Our HOD model was widely based on zheng2007; zehavi2011 and the halo occupation was inferred as a function of the halo mass. The HOD constraints in that work were important in various aspects of understanding the DES Y3 cosmology results, but are not suited for understanding the general galaxy population given that the lens samples had various selections that were imposed to optimize their photometric redshifts. In this work, we expand upon Zacharegkas2022 in three main ways, which we discuss in what follows. First, we incorporate theory predictions for galaxy clustering on small scales in addition to galaxy-galaxy lensing. In addition, our new model accounts for the stellar mass-information of the lens galaxies. This is done by direct parametrization of the HOD, but is equivalent to using, for example, the conditional stellar mass function mentioned before. Finally, we construct and validate a new, more generic stellar-mass-selected lens sample compared to what was used in Zacharegkas2022. This allows us to constrain the stellar-to-halo-mass relation of DES galaxies. We also note here that, the new stellar mass-selected galaxy sample can be used in future analyses of other science cases.
@@
[…]
@@ 2 The Halo Occupation Distribution model
In this section we discuss how we model the Halo Occupation Distribution (HOD) for this work. This model is based on Zacharegkas2022 and it is modified so that the stellar-mass information of the galaxies is accounted for. This is done by incorporating a stellar-to-halo-mass relation (SHMR) into the HOD framework. Below we introduce the SHMR used in this study (Section 2.1) and we then proceed to describe our HOD model (Section 2.2).
@@ 2.1 The stellar-to-halo-mass relation
Henceforth, we will be using the notation $f_{\rm SHMR}(M_{h})\equiv M_{\star}(M_{h})$ for the SHMR. There are a large number of variants of this relation used in the literature and, for the most part, they are in good agreement. However, especially at large masses, there can exist significant differences. In this work we follow the SHMR form in Kravtsov2018, which is the same as in Behroozi2013. In that paper the authors carefully calibrated the SHMR of high-mass halos and found that the relation in that regime is shallower than the HOD constraints by Leauthaud2012b and the abundance matching analyses by Moster2013; Behroozi2013. The authors suggest that the main reason is that in these previous studies the measured stellar-mass of the central galaxies was underestimated significantly.
@@
In the above expressions, we allow $\{M_{1},\epsilon,\alpha,\gamma,\delta\}$ to vary during our fits to the data. This functional form is a power-law with slope $\alpha$ at the low-mass limit, $M_{h}\ll M_{1}$, and breaks into a different power-law with slope $\gamma$ for high-mass halos, $M_{h}\gg M_{1}$, at the characteristic halo-mass scale $M_{1}$ with a characteristic stellar-mass-to-halo-mass ratio $\epsilon$. Finally, $\delta$ describes the strength of the power-law in the massive end of the relation.
@@ 2.2 The HOD parametrization
For a galaxy sample selected by a stellar-mass threshold $M_{\star}^{t}$, i.e. $M_{\star}\geq M_{\star}^{t}$, we can write down the HOD for the central galaxy similar to the form presented in zheng2007:
@@
To derive the HOD for a central occupation conditioned on a range of stellar masses $M_{\star,1},M_{\star,2}]$, we simply subtract two thresholded HOD’s from Equation (2) at $M_{\star,1}$ and $M_{\star,2}$:
@@
where $f_{\rm sat}$, $M_{\rm sat}$, $M_{\rm cut}$, $\alpha_{\rm sat}$ and $\kappa$ are free parameters, and we have defined $M_{\min}\equiv f_{\rm SHMR}^{-1}(M_{\star}^{t})$. Specifically, $f_{\rm sat}$ is the equivalent of $f_{\rm cen}$ in the central HOD of Equation (2) for the satellite occupation, $\alpha_{\rm sat}$ controls the power-law behaviour of the satellite occupation at large halo masses, and the stellar-mass dependence enters into the satellite modeling via the parameters $M_{\rm sat}$ and $M_{\rm cut}$ (see, e.g. leauthaud2011; zu2015), that we model as:
@@
where the new free parameters ($B_{\rm sat}$, $B_{\rm cut}$, $\beta_{\rm sat}$, $\beta_{\rm cut}$) have been introduced to allow the satellites to have a different form of HOD from the centrals. As an additional note, as opposed to Zacharegkas2022, in this current paper we decouple the HOD of the satellites in Equation (2.2) from the central one to give the model additional freedom in order to fit the more complex data we analyze here.
@@
The above approach effectively assumes a tophat distribution of the stellar mass in the range, and has been shown to be a good model for certain datasets (see e.g. leauthaud2011). This is also what we will use as our fiducial model in this work. We describe a more general case in Appendix A, where the stellar mass distribution takes on a more flexible form.
@@
$\qquad\times\left[2\langle N_{c}(M_{h}|M_{\star}^{t})\rangle+\lambda_{s}(M_{h})\rangle\right]\;.$ (9)
@@ 3 Modeling the observables
Equipped with the new HOD model, the main goal in this section is to combine the HOD model with the Halo Model (Seljak2000; Ma2000; PeacockSmith2000) to predict our two observables: galaxy-galaxy lensing (Section 3.1) and galaxy clustering (Section 3.2).
@@
as a function of distance $r$. Following hayashi2008; zu2015; clampitt2017, we combine the 1-halo and 2-halo term in the 3-dimensional real space, as it is found to more accurately describe observations at the quasi-linear regime. In the above, the 1-halo terms can be further decomposed into the central and satellite terms, as introduced in KravtsovHOD2004. We describe how we model each of these terms in $P_{\rm gm}$ and $P_{\rm gg}$ in Sections 3.1 and 3.2, respectively. We also note that each of the terms is a function of redshift ($z$) and scale ($k$). The central and satellite terms refer to the same definitions used in the HOD model in Section 2.2. The 1- and 2-halo terminology refers to whether the two tracers of interest (galaxy or dark matter) live in the same halo or two separate halos. There is a delicate choice to make when combining the 1- and 2-halo terms, which we discuss in Section 3.3. Once each term in Equations (10) and (11) are specified, together with the prescription for combining the two terms, it is projected into 2D (the projected spetra $C_{\rm gm}$ and $C_{\rm gg}$, of $P_{\rm gm}$ and $P_{\rm gg}$ respectively), then Fourier transformed into the real-space estimators that we measure ($\gamma_{t}$ and $w$) in Section 3.4. We describe the approach we take for incorporating the redshift and stellar mass uncertainties in Section 3.5, and finally we discuss any additional modeling components in Section 3.6